Anzio, Inc., has two classes of shares. Class B has10 times the voting rights as Class A. If you own 11% of the Class A shares and 25% of the Class B​ shares, what percentage of the total voting rights do you​ hold?

The percentage of the total voting rights that you hold is? (Round to one decimal​ place.)

Please show steps taken to get answer.
Answer #1

Let us assume Class A and Class shares are same in count.

Assume Class A shares are 100 and Class B shares are 100

Assume Voting Rights for Class A share = 1

Total Voting Rights for Class A Shares = 1 * 100 = 100

Total Voting Rights for Class A Shares = 10 * 100 = 1000

Total Voting Rights = 100 + 1000 = 1100

11% of Class A shares voting rights = 100 * 11%

= 11 rights

25% of Class B shares voting rights = 1000 * 25%

= 250 Rights

Total rights owned = 11 + 250

= 261

SHare = Total Rights owned / Total Rights

= 261 / 1100

= 0.2372 i.e 23.72%